Circular No. 6477

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1996 S2 (SHOEMAKER-HOLT 1) J. V. Scotti, University of Arizona, reports his recovery of comet

P/1987 U1 (= 1987z = 1988 VII) with the Spacewatch telescope at Kitt Peak, as given below. The indicated correction to the prediction by S. Nakano on MPC 23484 (ephemeris on MPC 26770) is ΔT = −0.89 day.

1996 UT  α_{2000}  δ_{2000}  m_{1}
Sept.19.19091  22^{h}51^{m}57.^{s}47  −2^{o}30′23′′.8  20.7
19.21169  22 51 56.81  −2 30 29.5  21.1
19.22904  22 51 56.20  −2 30 33.9  20.9
20.19518  22 51 23.97  −2 34 47.7
20.21330  22 51 23.39  −2 34 52.4
20.23085  22 51 22.72  −2 34 57.4  20.8
21.18698  22 50 51.25  −2 39 07.6
21.19863  22 50 50.85  −2 39 10.4  20.8
21.21013  22 50 50.47  −2 39 13.4  20.8

Improved orbital elements by G. V. Williams, Smithsonian Astrophysical Observatory, from 63 observations, 1987–1996 (mean residual 0′′.91):

Epoch = 1997 Nov. 8.0 TT

T = 1997 Nov. 20.2541 TT  ω = 210.^{o}2130
e = 0.321273  Ω = 214.5275 }2000.0
q = 3.047026 AU  i = 4.3617

a = 4.489325 AU n^{o} = 0.1036174 P = 9.512 years

1996TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{2}
Sept. 14 22^{h}54.^{m}95 − 2^{o}07.^{′}5 2.790 3.787 171.^{o}5 2.^{o}3 21.4
24 22 49.37 − 2 51.2 2.798 3.760 160.9  5.0 21.5
Oct. 4 22 44.62 − 3 32.1 2.834 3.733 149.9  7.7 21.7
14 22 41.13 − 4 07.2 2.895 3.707 139.1 10.1 21.8
24 22 39.21 − 4 33.9 2.977 3.680 128.6 12.2 21.9
Nov. 3 22 39.01 − 4 50.7 3.075 3.654 118.5 13.8 22.0

A more diffuse secondary condensation, 0.5 mag brighter than the primary on Sept. 20 UT, is separated by 0′.13 in p.a. 234^{o}, which corresponds to a current adjustment of the above orbit by ΔT ≃ +0.01 day. Coma diameters and m_{2} (Sept. 21.187) and tail lengths (Sept. 20): primary, 4′′, 21.6, 0′.09 in p.a. 241^{o}; secondary, 6′′, 21.9, 0′.14 in p.a. 246^{o}. The primary tail appears to bridge the gap between the two nuclei.

1996 September 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green